OpenStudy (anonymous):

Find the coordinates for the midpoint of the segment with endpoints given. (12, 4) and (-8, 8)

OpenStudy (nurali):

\[\frac{ x1+x2 }{ 2 },\frac{ y1+y2 }{ 2 }\] x1=12 x2=-8 y1=4 y2=8